# Detecting dilute axion stars constrained by fast radio bursts in the solar system via stimulated decay

### v1

@@ IV Abundance of dilute axion stars
This equation, applicable to axions produced via the misalignment mechanism, defines the lower bound of the axion parameter space, as shown in Fig. 1. By combining Eq. (4) with Eq. (13), we obtain the axion mass $m_{\phi}\simeq 1.17\times 10^{-6}$ eV and the decay constant $f_{a}\simeq 5.11\times 10^{12}$ GeV, under the assumption that QCD axions constitute the majority of dark matter. Substituting these values into Eq. (6), the maximum mass of the corresponding dilute axion star is calculated to be approximately $2.61\times 10^{-10}M_{\odot}$. This value coincides with the upper limit of the maximum axion star mass constrained by the parameter space derived from FRBs, as shown in Fig. 1. From Fig. 1, it is evident that the cosmological abundance line intersects the parameter space constrained by FRBs. This overlap significantly enhances the likelihood of axions existing within this region and provides additional support for the misalignment mechanism as a potential source of axions.
In the standard post-inflationary scenario, where $U(1)_{\rm{PQ}}$ symmetry is spontaneously broken after inflation, axion overdensities collapse during the radiation-dominated epoch, forming an early population of miniclusters with masses up to $10^{-12}M_{\odot}$ Hogan:1988mp ; Kolb:1993zz ; Kolb:1995bu . By redshift $z=100$, about $75\%$ of axion dark matter resides within these bound structures Eggemeier:2019khm , with minicluster masses reaching up to $10^{-9}M_{\odot}$ due to mergers. This closely matches the mass range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$ for axion stars constrained by FRBs. Dilute axion stars evolve toward a critical state through the accretion of axions from the surrounding environment or via merger events. Once they surpass this critical threshold, they collapse, emitting strong radio signals Di:2023nnb or relativistic axions Levkov:2016rkk . After the collapse, a residual dilute axion star remains, retaining a mass close to its critical value and entering a subcritical state. This suggests that the likelihood of dilute axion stars being in either a critical or subcritical state is relatively high. For simplicity, we will assume that dilute axion stars are in a critical state in the following discussion.
@@ V Echo from Dilute Axion Stars
A dilute axion star remains stable because the photon number resulting from spontaneous decay is insufficient to initiate stimulated decay. Now, consider transmitting radio waves with power $P$ to a dilute axion star. The coupled differential equations describing the evolution of the number of axions and photons within the dilute axion star are then modified as follows:
@@
where $N_{\gamma 0}\simeq 2PR_{\rm{AS}}/m_{\phi}$ represents the number of radio wave photons emitted from Earth and transmitted into the dilute axion star. By substituting Eq. (7) with $\kappa=1$ into $N_{\gamma 0}$, we obtain
$N_{\gamma 0}$ $\simeq$ $5.02\times 10^{28}\left(\frac{P}{50~{}\rm{MW}}\right)$ (17)
@@
Stimulated decay dominates over spontaneous decay when the condition $1{6\pi}(N_{\gamma}+N_{\gamma 0})/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})>$ 1 is satisfied. By transmitting radio waves with a power of $P={50~{}\rm{MW}}$ into the dilute axion star, we calculate $N_{\gamma 0}=3.86\times 10^{26}$ and find that ${6\pi}N_{\gamma 0}/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})=2.64\times 10^{15}\gg 1$, using the parameters $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV. Thus, radio waves of sufficient power, such as $P=50~{}\rm{MW}$, can effectively induce the stimulated decay of a dilute axion star. The multi-beam relativistic klystron amplifier currently has the capability to generate radio waves with an output power of 1.047 GW sun:2023 , significantly exceeding the required 50 MW. This substantial power output makes it a promising candidate for emitting radio waves to interact with axion stars. In the following discussion, we will use the example of a 50 MW radio wave received by an axion star.
Figure 2: Numerical evolution of the axion, photon, and “sterile” axion numbers within a dilute axion star under a powerful microwave radiation beam of $50{\rm{MW}}$. The solid line represents an axion star with a mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ and parameters $m_{\phi}=1.17\times 10^{-6}~{}{\rm{eV}}$ and $f_{a}=5.11\times 10^{12}~{}{\rm{GeV}}$. The dashed line corresponds to an axion star with a mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$ and parameters $m_{\phi}=5\times 10^{-5}~{}{\rm{eV}}$ and $f_{a}=5.20\times 10^{12}~{}{\rm{GeV}}$. Timing begins when the axion star is irradiated with radio waves. For $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, with a photon number of $N_{\gamma}=1.10\times 10^{21}$.
Figure 2 presents two numerical solutions showing the evolution of axion, photon, and “sterile” axion numbers for two representative parameter sets. The solid line corresponds to an axion star with a critical mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ with parameters $m_{\phi}=1.17\times 10^{-6}$ eV and a decay constant $f_{a}=5.11\times 10^{12}$ GeV, representing the intersection of cosmological abundance and QCD axion constraints in Fig. 1. The dashed line corresponds to an axion star with a critical mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, with parameters lying within the region constrained by FRBs: $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV. For the case $M_{\rm{AS}}=6.21\times 10^{-12}~{}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, resulting in a photon number $N_{\gamma}=1.10\times 10^{21}$. Considering the photons escaping from the dilute axion star, the luminosity can be expressed as:
@@
A notable feature of stimulated decay is that, in the axion rest frame, the emitted photons are restricted to propagate either in the same direction as or directly opposite to the incoming photons. However, the velocity distribution of axions within the dilute axion star introduces spatial dispersion of the emitted photons. Consequently, the luminosity flux observed from Earth is roughly given by
@@
where $d$ represents the distance between the dilute axion star and Earth. For a dilute axion star in its critical state, substituting Eq. (7) with $\kappa=1$ into $\Gamma_{e}=R_{\rm{AS}}^{-1}$, the photon escape rate becomes
$\Gamma_{e}\sim 1.24\times 10^{3}~{}{\rm{s^{-1}}}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right).$ (20)
By substituting Eq. (20) into Eq. (19), the flux can be expressed as
$F_{\phi}$ $=$ $3.50\times 10^{-35}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{2}\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right)$ (21)
@@
As illustrated in Fig. 1, the critical mass of dilute axion stars constrained by FRBs ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$. Substituting Eq. (6) with $\kappa=1$ into Eq. (21) for a mass of $6.21\times 10^{-12}M_{\odot}$, the luminosity flux of the dilute axion star is calculated as
@@
corresponding to the blue dot-dashed line with a flux of $F_{\phi}\approx 4.86\times 10^{-32}~{}{\rm{W/cm^{2}}}$ for $d=1000$ AU, as shown in Fig. 3. Similarly, substituting Eq. (6) with $\kappa=1$ into Eq. (21) for a mass of $2.61\times 10^{-10}M_{\odot}$ gives
@@
Figure 3: Flux of the dilute axion star at a distance of 1000 AU and the minimum detectable flux of SKA, FAST, ngLOBO, and LOFAR with an observation time of 1 hour. The sensitivity of these telescopes is detailed in Ref. Di:2024snm and its references. Even for LOFAR, the flux of the echoes is one to two orders of magnitude higher than its sensitivity, making the signal easily detectable and distinguishable.
Consequently, the stimulated decay of a dilute axion star within a volume of radius 1000 AU centered around Earth can be triggered by emitting a 50 MW radio beam into the star. The spectral line of the decay signal is nearly monochromatic, with a frequency of $f\simeq m_{\phi}/(4\pi)\approx 1.21(m_{\phi}/10^{-5}{\rm{eV}})~{}{\rm{GHz}}$, making it a distinct characteristic signal. Detection of such an echo would confirm the existence of axion stars, solve the dark matter puzzle, and provide crucial evidence linking some FRBs to axion stars. Furthermore, FRBs originating from collapsing axion stars could serve as standard candles to help resolve the Hubble tension Di:2024tlz . If such echoes are not detected within a 1000 AU volume around Earth using this method, constraints can be placed on the fraction of dark matter composed of dilute axion stars with critical masses constrained by FRBs. This fraction could be roughly less than $5\%$ for a mass of $6.21\times 10^{-12}M_{\odot}$ or $0.1\%$ for a mass of $2.61\times 10^{-10}M_{\odot}$, depending on the axion parameters. This estimate assumes a uniform distribution of axion stars throughout the Universe.
@@ VI Conclusions
Axions are a promising candidate for dark matter. Due to their bosonic nature, axions can achieve very high phase space densities, resulting in BEC and the formation of gravitationally bound structures known as axion stars. Collapsing axion stars with specific parameters may emit millisecond-long radio bursts with peak luminosities of $1.60\times 10^{42}$ erg/s, aligning with the characteristics of observed non-repeating FRBs. If some FRBs are indeed produced by collapsing axion stars, this would impose constraints on the axion parameter space, offering valuable guidance for experimental axion searches. Thus, exploring the axion parameter space in relation to FRBs is crucial, as it relates both to the origin of FRBs and the nature of dark matter.
@@
VII Acknowledgments
H. Di would like to thank Lijing Shao for useful discussions. This work was supported by National Natural Science Foundation of China under Grant No. 11947031 and East China University of Technology Research Foundation for Advanced Talents under Grant No. DHBK2019206.

### v4

@@ IV Abundance of dilute axion stars
This equation, applicable to axions produced via the misalignment mechanism, defines the lower bound of the axion parameter space, as shown in Fig. 1. By combining Eq. (4) with Eq. (13), we obtain the axion mass $m_{\phi}\simeq 1.17\times 10^{-6}$ eV and the decay constant $f_{a}\simeq 5.11\times 10^{12}$ GeV, assuming that QCD axions constitute the majority of dark matter. Substituting these values into Eq. (6) and taking $\kappa=1$, we calculate the maximum mass of the corresponding dilute axion star to be approximately $2.61\times 10^{-10}M_{\odot}$. This value coincides with the upper limit of the maximum axion star mass, as constrained by the parameter space derived from FRBs. As shown in Fig. 1, the cosmological abundance line intersects this parameter space. This overlap significantly increases the likelihood of axions existing in this region and further supports the misalignment mechanism as a potential source of axions.
In the standard post-inflationary scenario, where $U(1)_{\rm{PQ}}$ symmetry is spontaneously broken after inflation, axion overdensities collapse during the radiation-dominated epoch, forming an early population of miniclusters with masses up to $10^{-12}M_{\odot}$ Hogan:1988mp ; Kolb:1993zz ; Kolb:1995bu . By redshift $z=100$, about $75\%$ of axion dark matter resides within these bound structures Eggemeier:2019khm , with minicluster masses reaching up to $10^{-9}M_{\odot}$ due to mergers. This closely matches the mass range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$ for critical axion stars constrained by FRBs. Dilute axion stars evolve toward a critical state through the accretion of axions from the surrounding environment or via merger events. Once they surpass this critical threshold, they collapse, emitting strong radio signals Di:2023nnb or relativistic axions Levkov:2016rkk . After the collapse, a residual dilute axion star remains, retaining a mass close to its critical value and entering a subcritical state. This suggests that the likelihood of dilute axion stars being in either a critical or subcritical state is relatively high. For simplicity, we will assume that dilute axion stars are in a critical state in the following discussion.
@@ V Echo from Dilute Axion Stars
A dilute axion star remains stable because the photon number resulting from spontaneous decay is insufficient to initiate stimulated decay. However, by transmitting a powerful radio beam with power $P$ to the axion star, with an angular frequency of photons equal to half the axion mass, the stimulated decay of axions can be significantly enhanced. As a result, the coupled differential equations given by Eq. (10) and Eq. (11) are modified as follows:
@@
where $N_{\gamma 0}\simeq 2PR_{\rm{AS}}/m_{\phi}$ represents the number of photons from the radio beam transmitted into the axion star. By substituting Eq. (7) into $N_{\gamma 0}$, we obtain
$N_{\gamma 0}$ $\simeq$ $5.02\times 10^{28}~{}\kappa^{1/2}\left(\frac{P}{50~{}\rm{MW}}\right)$ (17)
@@
Stimulated decay dominates over spontaneous decay when the condition $1{6\pi}(N_{\gamma}+N_{\gamma 0})/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})>$ 1 is satisfied. By transmitting a powerful radio beam with a power of $50~{}\rm{MW}$ into the axion star, we obtain $N_{\gamma 0}=3.86\times 10^{26}$ and find that ${6\pi}N_{\gamma 0}/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})=2.64\times 10^{15}\gg 1$, using the parameters $m_{\phi}=5\times 10^{-5}$ eV, $f_{a}=5.20\times 10^{12}$ GeV and taking the model-dependent constant $\kappa=1$. Thus, radio waves of sufficient power, such as $P=50~{}\rm{MW}$, can effectively induce the stimulated decay of a dilute axion star. The multi-beam relativistic klystron amplifier currently has the capability to output radio waves with a power of 1.047 GW sun:2023 , significantly exceeding the required 50 MW. This substantial power output makes it a promising candidate for emitting radio waves to interact with axion stars. In the following discussion, we will use the example of a 50 MW radio wave received by an axion star.
Figure 2: Evolving axion, photon, and “sterile” axion numbers inside a dilute axion star under a $50{\rm{MW}}$ radio beam. The solid line represents an axion star with a mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ and parameters $m_{\phi}=1.17\times 10^{-6}~{}{\rm{eV}}$ and $f_{a}=5.11\times 10^{12}~{}{\rm{GeV}}$. The dashed line corresponds to an axion star with a mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$ and parameters $m_{\phi}=5\times 10^{-5}~{}{\rm{eV}}$ and $f_{a}=5.20\times 10^{12}~{}{\rm{GeV}}$. Timing begins when the axion star is irradiated with radio beam. For $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, with a photon number of $N_{\gamma}=1.10\times 10^{21}$.
Figure 2 presents two numerical solutions illustrating the evolution of the axion, photon, and “sterile” axion numbers for two representative parameter sets, assuming the model-dependent constants $K=1$ and $\kappa=1$. The solid line represents an axion star with a critical mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ with parameters $m_{\phi}=1.17\times 10^{-6}$ eV and a decay constant $f_{a}=5.11\times 10^{12}$ GeV, representing the intersection of cosmological abundance and QCD axion constraints in Fig. 1. The dashed line represents an axion star with a critical mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, with parameters lying within the region constrained by FRBs: $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV. For the case $M_{\rm{AS}}=6.21\times 10^{-12}~{}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, resulting in a photon number $N_{\gamma}=1.10\times 10^{21}$. Considering the photons escaping from the dilute axion star, the luminosity can be expressed as:
@@
A notable feature of axion stimulated decay is that, in the axion rest frame, the emitted photons are restricted to propagate either in the same direction as or directly opposite to the incoming photons. However, within the dilute axion star, the velocity distribution of the axions leads to spatial dispersion of the emitted photons. Consequently, the luminosity flux observed from Earth is roughly given by
@@
where $d$ denotes the distance between the dilute axion star and Earth. For a dilute axion star in its critical state, substituting Eq. (7) into $\Gamma_{e}=R_{\rm{AS}}^{-1}$, the photon escape rate becomes
$\Gamma_{e}\sim 1.24\times 10^{3}~{}{\rm{s^{-1}}}~{}\kappa^{-1/2}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right).$
By substituting Eq. (V) into Eq. (19), the flux can be expressed as
$F_{\phi}$ $=$ $3.50\times 10^{-35}~{}\kappa^{-1/2}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{2}\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right)$ (21)
@@
As illustrated in Fig. 1, the critical mass of dilute axion stars constrained by FRBs ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$. Substituting Eq. (6) into Eq. (21) for a mass of $6.21\times 10^{-12}M_{\odot}$, the luminosity flux of the dilute axion star is calculated as
@@
corresponding to the blue dot-dashed line with a flux of $F_{\phi}\approx 4.86\times 10^{-32}~{}{\rm{W/cm^{2}}}$ for $d=1000$ AU, as shown in Fig. 3. Similarly, substituting Eq. (6) into Eq. (21) for a mass of $2.61\times 10^{-10}M_{\odot}$ gives
@@
Figure 3: Axion star flux at 1000 AU and the sensitivity of SKA, FAST, ngLOBO, and LOFAR for 1 hour observations. The sensitivity of these telescopes is detailed in Ref. Di:2024snm and its references. Even for LOFAR, the flux of the echoes is one to two orders of magnitude higher than its sensitivity, making the signal easily detectable and distinguishable.
Consequently, the stimulated decay of a dilute axion star within a volume of radius 1000 AU centered around Earth can be triggered by emitting a 50 MW radio beam into the star. The decay signal produces a spectral line that is almost monochromatic, with a frequency given by $f\simeq m_{\phi}/(4\pi)\approx 1.21(m_{\phi}/10^{-5}{\rm{eV}})~{}{\rm{GHz}}$, which serves as a distinctive feature of the signal. Detection of such an echo would confirm the existence of axion stars, solve the dark matter puzzle, and provide crucial evidence linking some FRBs to axion stars. Furthermore, FRBs originating from collapsing axion stars could serve as standard candles to help resolve the Hubble tension Di:2024tlz . If such echoes are not detected within a 1000 AU volume around Earth using this method, constraints can be placed on the fraction of dark matter composed of dilute axion stars with critical masses constrained by FRBs. This fraction could be roughly less than $5\%$ for a mass of $6.21\times 10^{-12}M_{\odot}$ or $0.1\%$ for a mass of $2.61\times 10^{-10}M_{\odot}$, depending on the axion parameters. This estimate assumes a uniform distribution of axion stars throughout the Universe.
@@ VI Conclusions
Axions are a promising candidate for dark matter. Due to their bosonic properties, they can reach exceptionally high phase space densities, which may lead to BEC and the formation of gravitationally bound objects known as axion stars. Collapsing axion stars with specific parameters may emit millisecond-long radio bursts with peak luminosities of approximately $10^{42}$ erg/s, which are consistent with the characteristics of observed non-repeating FRBs. If some FRBs are indeed produced by collapsing axion stars, this would impose constraints on the axion parameter space, offering valuable guidance for experimental axion searches. Thus, exploring the axion parameter space in relation to FRBs is crucial, as it relates both to the origin of FRBs and the nature of dark matter.
@@
Acknowledgements.
H. D. would like to thank Lijing Shao for useful discussions. This work was supported by National Natural Science Foundation of China under Grant No. 11947031 and East China University of Technology Research Foundation for Advanced Talents under Grant No. DHBK2019206.